# Efficient Generation of Smooth Paths with Curvature Guarantees by Mollification

### v1

@@ V-B The general case
We propose the following methodology. Suppose we have $p\in\mathbb{N}$ segments with $p\geq 2$. Using equation (8) and given a maximum curvature $\kappa_{\max}>0$, we can compute for each pair of consecutive segments its respective [^3] The index $i$ ranges from the first pair to the last pair of consecutive segments. $\varepsilon_{i}>0$ such that, under the three-point-two-segment approximation, their curvatures are upper bounded. If $\varepsilon_{i}<\frac{1}{2}$ for all $i$, then equation (8) is exact, because only the two segments used for computing $\varepsilon_{i}$ contribute to the mollification at the junction point. In this case, take $\varepsilon=\max_{i}\varepsilon_{i}$, which is valid because the only dependence on $\varepsilon_{i}$ in the right-hand side of equation (8) is through $\frac{1}{\varepsilon_{i}}$; hence, $\varepsilon$ satisfies the bound for each pair of segments. If $\varepsilon_{i}>\frac{1}{2}$ for some $i$, then equation (8) becomes an approximation. In this case, one can either accept an admissible error or use $\varepsilon=\max_{i}\varepsilon_{i}$ as an initial condition for an optimization algorithm that seeks the minimum $\varepsilon>0$ that upper bounds the curvature. In either case, equation (8) is a powerful, computationally inexpensive tool that can be used to either compute an exact upper bound for the complete trajectory or reduce computation time in an optimization algorithm.
VI Numerical validations and real experiments
@@
To demonstrate the effectiveness of our path generation approach, we present both numerical and experimental results for path following of a mollified path by a unicycle vehicle. Specifically, we employ the Singularity-Free Guiding Vector Fields (SF-GVF) path following algorithm [17, 14]. In brief, SF-GVF takes a parametric path $f\in C^{2}(\mathbb{R},\mathbb{R}^{n})$ as input and constructs a vector field $\chi\in C^{2}(\mathbb{R}^{n},\mathbb{R}^{n})$ whose flow traces the mollified path.
VI-A Numerical results
@@
We are going to consider the so called “heart” function as our input path. Define the function
@@
Figure 8: Representation of the numerical simulation. In both pictures the original path $f$, is shown as a blue dotted line, while as a red dotted line the mollified trajectory $F$ for $\varepsilon_{1}=\varepsilon_{2}=0.4$ in the left picture, and for $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$ in the right picture. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
@@
A numerical simulation of the vehicle under SF-GVF using $\varepsilon_{1}=\varepsilon_{2}=0.4$ for the “heart” path is shown in the left plot of Figure 8. The vehicle’s trajectory indicates convergence to the desired mollified path. Moreover, as $\varepsilon\to 0$, the trajectory approaches the original path more closely. Note that the mollified path lies inside the original path, as predicted by Theorem 6. In practical terms, this means that the “heart” function can now be used with SF-GVF, extending the applicability of this path following algorithm. Clearly, to improve convergence to the original path, we can reduce both $\varepsilon_{1}$ and $\varepsilon_{2}$, since by Theorem 1 we have uniform convergence on compact sets as $\varepsilon_{1},\varepsilon_{2}\to 0$. However, to demonstrate the flexibility of the approach, we also consider the case $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$, whose simulation is presented in the right plot of Figure 8. Note that in the first component, the mollified curve is better adjusted to the original curve, while in the second component, the opposite occurs. This results from $\varepsilon_{2}$ being four times larger than $\varepsilon_{1}$. Indeed, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ can be constrained by the vehicle’s dynamics. This is a key advantage of the method: by simply adjusting these parameters, we can ensure that the vehicle follows the curve within its dynamic limits, thereby avoiding issues related to reconverging to the path. Moreover, numerical computations show that the length of the original path with respect to the $\ell_{1}$ norm is (in arbitrary units) $25.58$, while the mollified curve has length $23.16$ in the left case and $21.74$ in the right case of Figure 8, as predicted by Theorem 8. The same conclusions from Theorem 8 can be verified with respect to any other arbitrary $\ell_{p}$ norm.
@@
Figure 9: Representation of the three dimensional numerical simulation. The original path $f$ is represented as a solid blue line, while the mollified trajectory $F=(F_{k})_{k=1}^{3}=(f_{k}*\varphi_{\varepsilon_{k}})_{k=1}^{3}$ is shown as a dashed red line, with $\varphi$ as in example 1 and $\varepsilon_{k}=1$ for $k\in\{1,2,3\}$. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [14] starting from an arbitrary initial position.
@@
Finally, for completeness, a numerical simulation of a 3D mollified path is presented in Figure 9. The notation used is identical to that in the previous numerical simulations. The original path $f$ is constructed via linear interpolation between a sequence of vertices/waypoints of a three-dimensional cube. As can be seen, $f$ is non-differentiable at these vertices. In contrast, the mollified function $F$ provides a smooth approximation that can be effectively employed in SF-GVF, as illustrated in Figure 9. Indeed, the flow of the guiding vector field converges to the mollified trajectory. Moreover, Theorem 8 can also be validated numerically. In this case, the length of the original path in the $\ell_{2}$ norm is (in arbitrary units) $7.38$, while the length of the mollified path is $5.48$. An important feature of this approach is its scalability: any trajectory of the form $g:\mathbb{R}\to\mathbb{R}^{n}$ can be mollified component-wise, producing a sufficiently smooth curve that can be further adapted to satisfy a variety of constraints.
VI-B Experimental results
@@
Before presenting the experimental results, we introduce the software and hardware platforms used in the experiments. We also provide the necessary links to the developed software so any interested reader can replicate these experiments.
VI-B1 Software platform
We use Paparazzi UAV [18] as a development environment for programming autonomous vehicles. Paparazzi is a free and open-source hardware and software project designed to create a flexible autopilot system. Although it was initially developed for unmanned aerial vehicles (UAVs), researchers from various universities have extended its applications to other autonomous vehicles. Moreover, Paparazzi supports multiple hardware designs, making the interconnection between different modules accessible; the user simply needs to focus on the desired algorithm to implement, e.g., mollification of trajectories. Another important component of Paparazzi is its Ground Control Station (GCS), where vehicle positions are visualized and different sensor and internal variables can be monitored via telemetry. We have added the necessary code to define, display, and modify the mollification of straight lines on the fly. See Figure 10 for a screenshot of the Paparazzi GCS showing the mollification of a linear interpolation of points in $\mathbb{R}^{2}$.
@@ VI-B1 Software platform
Using the Paparazzi GCS, the user can issue real-time high-level commands to the Rover’s autopilot via telemetry. Examples of such commands include path following instructions for desired waypoints, speed setpoints, and guidance and speed controller constants. The rest of the system runs on the vehicle’s onboard computer, enabling fully autonomous operation. It is important to note that waypoints can be adjusted to fit the trajectory into the desired area. By simply moving the control or endpoints, the user can shape the path as desired. Onboard software in a microcontroller recalculates the trajectory by mollifying it using the user-specified parameters whenever a point or parameter is changed. This provides considerable flexibility in experimental environments where positional uncertainties may arise.
VI-B2 Experimental platform
Figure 11: Rover vehicle and hardware used during the experiments.
The experimental hardware platform is presented in Figure 11 and can be modeled using the standard unicycle model. The rover is equipped with a Matek F765-Wing Autopilot as its central component. The autopilot consists of an STM32 microcontroller and integrated Inertial Measurement Units (IMUs) for inertial navigation, along with input/output ports for connecting various sensors including barometers, GNSS, compass, and radio receivers. We use the open-source Paparazzi UAV framework for programming and uploading code to the autopilot. Paparazzi supports the sensors required for outdoor autonomous operation: GNSS receivers and compass for accurate positioning and yaw attitude, radio receivers as a fail-safe for manual control during algorithmic failures, and telemetry radios enabling ground reception of vehicle states and transmission of commands to the vehicle. Our implementation uses the Matek M10Q-5883 GNSS and compass module, a Futaba 7008SB radio receiver, and a Zigbee Xbee telemetry radio.
VI-B3 Experimental results
@@
Experimental data extracted from Paparazzi logs is shown in Figure 12. We created a continuous but non-differentiable path by linearly interpolating points in $\mathbb{R}^{2}$ as it is done in Section V. As noted above, any Paparazzi user can create such trajectories by simply moving points in the ground control station before or during the experiment in real time. The experiment uses a curve similar to that in Figure 10.
The experimental objective is as follows: Given a desired non-differentiable path $f$, mollify it with parameter $\varepsilon>0$ to obtain $F_{\varepsilon}=(f*\varphi_{\varepsilon})$, where $\varepsilon$ is a function of vehicle speed limited by maximum allowed curvature. To demonstrate potential applications, we use a linear relationship between maximum allowed curvature and speed. For speed $v>0$, with $R_{min}$, $R_{max}$, and $v_{max}$ denoting minimum radius, maximum radius, and maximum speed respectively, the allowed radius of curvature is $R(v):=R_{min}+\frac{v}{v_{max}}(R_{max}-R_{min})$. Note this is merely illustrative; curvature and speed need not be linearly related in practice.
We adapt the curve to vehicle dynamics as follows: knowing the speed at each segment midpoint and assuming it remains constant until the next segment midpoint, we compute maximum allowed curvature using the linear relationship, then obtain the minimum allowed $\varepsilon>0$ from equation (8). This process applied to each segment pair creates a family of curves for different $\varepsilon>0$ values. With six segments, we obtain five mollified curves (one per corner) plus one for the default $\varepsilon>0$ value. As before, $f:\mathbb{R}\to\mathbb{R}^{2}$ denotes the original path, $\varphi$ is the mollifier from Example 1, $F_{\varepsilon}=(f*\varphi_{\varepsilon})$ is the speed-dependent mollified path, and $r:[0,\infty)\to\mathbb{R}^{2}$ denotes vehicle position as a function of time, with $r(t;r_{0})$ indicating initial position $r_{0}\in\mathbb{R}^{2}$.
Figure 12 shows the results. The family of six mollified paths adapts well to vehicle dynamics, approaching the original path when curvature constraints allow (e.g., segments four and five) while strongly constraining the curve when necessary (e.g., the last two segments). The theoretical results hold: mollified curves lie within the convex hull of the original path (Theorem 6), and parameters are sufficiently small that Propositions 2 and 5 apply. These experiments validate our approach, demonstrating both theoretical solution of Problem 1 and practical viability on real and affordable hardware.
@@
Figure 12: Experiments with the real rover vehicle. Left column figures represent the original trajectory in black, the mollified family trajectories as red dashed lines, the complete trajectory of the vehicle as a dashed blue line, its initial point as a blue cross, and the position of the vehicle for its corresponding $\varepsilon_{i}>0$ in a thick solid blue line. Right column figures represent in a similar fashion, the error between the rover’s position and the path in blue, and the GNSS accuracy in green.
VII Conclusions

### v2

@@ 5.2 The general case
We propose the following methodology. Suppose we have $p\in\mathbb{N}$ segments with $p\geq 2$. Using (8) and given a maximum curvature $\kappa_{\max}>0$, we can compute for each pair of consecutive segments its respective [^1] The index $i$ ranges from the first pair to the last pair of consecutive segments. $\varepsilon_{i}>0$ such that, under the three-point-two-segment approximation, their curvatures are upper bounded. If $\varepsilon_{i}<\frac{1}{2}$ for all $i$, then (8) is exact, because only the two segments used for computing $\varepsilon_{i}$ contribute to the mollification at the junction point. In this case, take $\varepsilon=\max_{i}\varepsilon_{i}$, which is valid because the only dependence on $\varepsilon_{i}$ in the right-hand side of (8) is through $\frac{1}{\varepsilon_{i}}$; hence, $\varepsilon$ satisfies the bound for each pair of segments. If $\varepsilon_{i}>\frac{1}{2}$ for some $i$, then (8) becomes an approximation. In this case, one can either accept an admissible error or use $\varepsilon=\max_{i}\varepsilon_{i}$ as an initial condition for an optimization algorithm that seeks the minimum $\varepsilon>0$ that upper bounds the curvature. In either case, (8) is a powerful, computationally inexpensive tool that can be used to either compute an exact upper bound for the complete trajectory or reduce computation time in an optimization algorithm.
6 Numerical validations and real experiments
@@
To demonstrate the effectiveness of our path generation approach, we first show a comparison between the path generated by different spline methods and the one by mollification. We also present both numerical and experimental results for path following of a mollified path by a unicycle vehicle. Specifically, we employ the Singularity-Free Guiding Vector Fields (SF-GVF) path following algorithm [18, 17]. In brief, SF-GVF takes a parametric path $f\in C^{2}(\mathbb{R},\mathbb{R}^{n})$ as input and constructs a vector field $\chi\in C^{2}(\mathbb{R}^{n},\mathbb{R}^{n})$ whose flow traces the mollified path.
6.1 Comparison with traditional interpolation methods
We recall that mollification can generate paths from more general inputs than those for which interpolation is possible, such as the input that considers only $x\in\mathbb{Q}$ in Remark 4.26. Nonetheless, to compare traditional interpolation methods with our path generation via mollification, consider a collection of points in $\mathbb{R}^{2}$ that are linearly interpolated, and their $C^{2}$ cubic splines, B-splines and quintic Hermite interpolation splines, i.e, polynomial splines that are twice continuously differentiable and which are computed to specifically pass through the pre-defined collection of points. We compare the generated paths between the spline approaches and the proposed mollification in Figure 8.
Figure 8: Comparison of different approach for path generation. The blue line represents the path generated from the linearly interpolated sequence of points shown in black. The mollifier used is the one presented in Example 1 with $\varepsilon=0.4$.
As it can be seen, the mollification approach is the method that resembles the original function the most. Moreover, due to Proposition 5.38 we know exactly in which intervals the original function and the mollified one coincide. Clearly, this is at the cost of not intersecting the collection of points, except the initial and ending ones. Nevertheless, note that numerically speaking, the mollification method is by far the simplest one. Indeed, for $C^{2}$ continuity there are several conditions that the coefficients of the spline polynomials must met, and which are extremely sensible to changes in the path. That is, a small change in the original path, i.e., a small change in a point of the collection, can generate a big change in the generated path. This is not something that happens in mollification, since by Proposition 5.38 and the results of Section 4 we know exactly in which sets the original path and its mollification coincide, where is the mollified curve enclosed, as well as the geometric properties preserved. Finally, we want to remark an important fact. The polynomials spline here presented are at most $C^{2}$ continuous at the collection of points, and its maximum curvature is not easy to compute between two segments nor the complete spline. For this purpose an optimization approach is often carried as mentioned in the introduction. In the mollified case, with just a single operation, $C^{\infty}$ continuity is obtained, and its derivatives are easy to compute thanks to Theorem 1. We have also provided, for this specific example, an analytical and simple to compute curvature upper bound in (8). This implies that mollified path can be computed or changed online in low-cost platforms.
6.2 Usage with a path following algorithm
@@
For the path-following part, we are going to consider the so called “heart” function as our input path. Define the function
@@
Figure 9: Representation of the numerical simulation. In both pictures the original path $f$, is shown as a blue dotted line, while as a red dotted line the mollified trajectory $F$ for $\varepsilon_{1}=\varepsilon_{2}=0.4$ in the left picture, and for $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$ in the right picture. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [17] starting from an arbitrary initial condition (IC).
@@
A numerical simulation of the vehicle under SF-GVF using $\varepsilon_{1}=\varepsilon_{2}=0.4$ for the “heart” path is shown in the left plot of Figure 9. The vehicle’s trajectory indicates convergence to the desired mollified path. Moreover, as $\varepsilon\to 0$, the trajectory approaches the original path more closely. Note that the mollified path lies inside the original path, as predicted by Theorem 4.28. In practical terms, this means that the “heart” function can now be used with SF-GVF, extending the applicability of this path following algorithm. Clearly, to improve convergence to the original path, we can reduce both $\varepsilon_{1}$ and $\varepsilon_{2}$, since by Theorem 1 we have uniform convergence on compact sets as $\varepsilon_{1},\varepsilon_{2}\to 0$. However, to demonstrate the flexibility of the approach, we also consider the case $\varepsilon_{1}=0.2$ and $\varepsilon_{2}=0.8$, whose simulation is presented in the right plot of Figure 9. Note that in the first component, the mollified curve is better adjusted to the original curve, while in the second component, the opposite occurs. This results from $\varepsilon_{2}$ being four times larger than $\varepsilon_{1}$. Indeed, the values of $\varepsilon_{1}$ and $\varepsilon_{2}$ can be constrained by the vehicle’s dynamics. This is a key advantage of the method: by simply adjusting these parameters, we can ensure that the vehicle follows the curve within its dynamic limits, thereby avoiding issues related to reconverging to the path. Moreover, numerical computations show that the length of the original path with respect to the $\ell_{1}$ norm is (in arbitrary units) $25.58$, while the mollified curve has length $23.16$ in the left case and $21.74$ in the right case of Figure 9, as predicted by Theorem 4.34. The same conclusions from Theorem 4.34 can be verified with respect to any other arbitrary $\ell_{p}$ norm.
@@
Figure 10: Representation of the three dimensional numerical simulation. The original path $f$ is represented as a solid blue line, while the mollified trajectory $F=(F_{k})_{k=1}^{3}=(f_{k}*\varphi_{\varepsilon_{k}})_{k=1}^{3}$ is shown as a dashed red line, with $\varphi$ as in example 1 and $\varepsilon_{k}=1$ for $k\in\{1,2,3\}$. The black solid line represents the (now smooth) flow generated by the guiding vector field according to [17] starting from an arbitrary initial condition (IC).
6.3 Three dimensional path mollification
@@
Finally, for completeness, a numerical simulation of a 3D mollified path is presented in Figure 10. The notation used is identical to that in the previous numerical simulations. The original path $f$ is constructed via linear interpolation between a sequence of vertices/waypoints of a three-dimensional cube. As can be seen, $f$ is non-differentiable at these vertices. In contrast, the mollified function $F$ provides a smooth approximation that can be effectively employed in SF-GVF, as illustrated in Figure 10. Indeed, the flow of the guiding vector field converges to the mollified trajectory. Moreover, Theorem 4.34 can also be validated numerically. In this case, the length of the original path in the $\ell_{2}$ norm is (in arbitrary units) $7.38$, while the length of the mollified path is $5.48$. An important feature of this approach is its scalability: any trajectory of the form $g:\mathbb{R}\to\mathbb{R}^{n}$ can be mollified component-wise, producing a sufficiently smooth curve that can be further adapted to satisfy a variety of constraints.
6.4 Experimental results
@@
Before presenting the experimental results, we introduce the software and hardware platforms used in the experiments. We also provide the necessary links to the developed software so that any interested reader can replicate these experiments.
6.4.1 Experimental and Software Platform
Our experimental platform, shown in Figure 11, is a rover modeled as a unicycle, built around a Matek F765-Wing autopilot with an STM32 microcontroller, integrated IMUs, and support for GNSS, compass, and radio receivers—specifically a Matek M10Q-5883, a Futaba 7008SB receiver, and a Zigbee Xbee telemetry radio. The entire system runs on the open-source Paparazzi UAV framework [6], which handles autonomous operation, telemetry, and real-time communication with the Ground Control Station (GCS). Through the GCS, the user can issue high-level commands and adjust waypoints on the fly, while the onboard microcontroller recomputes the mollified path in real time whenever a point or parameter is modified, as illustrated in Figure 12.
Figure 11: Rover vehicle and hardware used during the experiments.
@@ 6.4.1 Experimental and Software Platform
6.4.2 Experimental data
@@
We show our logs in Figure 13. We created a continuous but non-differentiable path by linearly interpolating points in $\mathbb{R}^{2}$ as it is done in Section 5. As noted above, any Paparazzi user can create such trajectories by simply moving points in the ground control station before or during the experiment in real time. The experiment uses a curve similar to that in Figure 12.
@@
Figure 13: Real experiment of path following of a mollified non-differentiable path with a rover vehicle. The original non-differentiable path $f$ is in black, the mollified family trajectories $\{F_{\varepsilon_{i}}\}_{i=1}^{6}$ to be followed at each stage are in solid red, its initial point $r_{0}$ is denoted as the blue cross, the position of the vehicle for its corresponding $\varepsilon_{i}>0$ at times $t\in[t_{i-1},t_{i}]$ in a thick solid blue line. Finally, the complete trajectory of the vehicle $r(t;r_{0})$ throughout all the stages is the dashed blue line.
The experimental objective is as follows: Given a desired non-differentiable path $f$, mollify it with parameter $\varepsilon>0$ to obtain $F_{\varepsilon}=(f*\varphi_{\varepsilon})$, where $\varepsilon$ is a function of vehicle speed limited by maximum allowed curvature. To demonstrate potential applications, we use a linear relationship between maximum allowed curvature and speed. For speed $v>0$, with $R_{min}$, $R_{max}$, and $v_{max}$ denoting minimum radius, maximum radius, and maximum speed respectively, the allowed radius of curvature is $R(v):=R_{min}+\frac{v}{v_{max}}(R_{max}-R_{min})$. Note this is merely illustrative; curvature and speed need not be linearly related in practice. We are going to consider a curve consisting on seven two dimensional points which are linearly interpolated, as shown in Figure 13.
The adaptation of the curve to vehicle dynamics operates as follows: at initial time $t_{0}$, we set $\varepsilon_{1}=0.5$ to generate the first mollified curve $F_{\varepsilon_{1}}=f*\varphi_{\varepsilon_{1}}$, where $f$ is the original curve and $\varphi$ is as in Example 1. As the vehicle advances and reaches each segment midpoint at times $\{t_{i}\}_{i=1}^{6}$, the speed is measured and used to compute the maximum allowed curvature, from which the minimum permissible $\varepsilon_{i+1}$ is determined via (8). This generates a family of six parameters $\{\varepsilon_{i}\}_{i=1}^{6}$ and corresponding mollified curves $\{F_{\varepsilon_{i}}\}_{i=1}^{6}$ that dynamically adapt to vehicle constraints. As shown in Figure 13, the resulting paths remain close to the original trajectory where curvature permits, e.g., segments four and five, while strongly constraining the curve when necessary, e.g., the last two segments. All theoretical guarantees hold: the mollified curves lie within the convex hull of the original path (Theorem 4.28), and parameters are sufficiently small that Propositions 4.6 and 4.16 apply. These experiments validate both the theoretical solution of Problem 1 and the practical viability of our approach on real, affordable hardware.
7 Conclusions